OpenStudy (anonymous):

Integrate : ∫dx/(sqrt((x^2)-16))

OpenStudy (anonymous):

If you remember your pythagorean identities, you can ALWAYS tell. sin^2(x) + cos^2(x) = 1 Divide by cos^2(x) tan^2(x) + 1 = sec^2(x) or tan^2(x) = sec^2(x) - 1 Divide by sin^2(x) 1 + cot^2(x) = csc^2(x) or cot^2(x) = csc^2(x) - 1 The one it looks most like is the one you pick. In this case, x^2 + 4^2 means we'll need this one tan^2(x) + 1 = sec^2(x). Multiply by 4^2.

OpenStudy (anonymous):

did you just copy and paste from this previous question lol http://openstudy.com/study#/updates/5085e59de4b0b7c30c8e3fa5

OpenStudy (anonymous):

i already checked it out still dont get it :(

OpenStudy (anonymous):

Yes I did lol But I know it works as this is a simple technique I mostly use in pythagorean identities

OpenStudy (anonymous):

First off, You must have your integral be something like this: 1/sqrt(1+u^2) where it can be anything so it is easier this way.

OpenStudy (anonymous):

yea but how do you get to that point, that all i wanna know

OpenStudy (anonymous):

Do you need the formula step-by-step or just the answer itself with an explanation?

OpenStudy (anonymous):

just how to get to something like 1/sqrt(1+u^2)

hartnn (hartnn):

when you have the form \(\Large \sqrt{x^2-a^2}\) in the denominator, put \(\Large x= a\sec t \) so, here, put x =4 sec t dx = ... ?

OpenStudy (anonymous):

4sec t tan t ?

hartnn (hartnn):

yes, and your denominator now becomes ?

OpenStudy (anonymous):

i dont get it bro, like should i substitute this in denominator ?

hartnn (hartnn):

yes, \(\sqrt{x^2-a^2 }= \sqrt {a^2\sec^2t-a^2} = ... ?\) ccan you simplify ?

OpenStudy (anonymous):

a sec t - a

hartnn (hartnn):

nopes...

hartnn (hartnn):

\(\sqrt{x^2-a^2 }= \sqrt {a^2\sec^2t-a^2} = \sqrt{a^2(sec^2t-1)}=a \sqrt {\tan^2 t}=a\tan t\) got this

hartnn (hartnn):

oh and here a=4

OpenStudy (anonymous):

i got it buy that i should put x=a sect ? , is it a formula ?

hartnn (hartnn):

we put x = a sec t just so that denominator would simplify to a tan t! like if we had , sqrt (a^2-x^2) we would have put x = a sin t , so that denominator would become a cos t!

OpenStudy (anonymous):

if it was (a^2-x^2) i would simple use arcsin(a/x) + c formula but anyways, now the denominator becomes 4 tan t , how does that helps

hartnn (hartnn):

notice that numerator is 4 sec t tant what gets cancelled ?

OpenStudy (anonymous):

we are left with sec t

hartnn (hartnn):

yes, just integral sec t dt can you integrate that ?

OpenStudy (anonymous):

ln |sec t + tan t| + C ?

hartnn (hartnn):

correct! now just plug back in t = sec inverse x

hartnn (hartnn):

sorry, t = sec inverse (x/4)

hartnn (hartnn):

because x = 4 sec t

hartnn (hartnn):

no... ln |x/4 + tan (sec inverse (x/4))| +c do you know how to simplify tan (sec inverse (x/4) ?

OpenStudy (anonymous):

plz explain how u got 4 sec t tant in the numerator and how i got cancelled

OpenStudy (anonymous):